# Non-perturbative self-consistent electron-phonon spectral functions and transport

### v2

@@ Abstract
Electron-phonon coupling often dominates the electron spectral functions and carrier transport properties. However, studies of this effect in real materials have largely relied on perturbative one-shot methods due to the lack of a first-principles theoretical and computational framework. Here, we present a self-consistent theory and implementation for the non-perturbative calculations of spectral functions and conductivity due to electron-phonon coupling. Applying this method to monolayer InSe, we demonstrate that self-consistency qualitatively affects the spectral function and transport properties compared to state-of-the-art one-shot calculations and allow one to reconcile experimental angle-resolved photoemission experiments. The developed method can be widely applied to materials with dominant electron-phonon coupling at moderate computational cost.
Spectral functions are essential in many-body physics, linking theory and experiment. Experimentally, spectral functions can be extracted using high-resolution angle-resolved photoemission spectroscopy [1]. Theoretically, spectral functions can be calculated from Green’s functions, including many-body correlation via the self-energy. The coupling between electrons and phonons, the collective vibration of the lattice, plays a crucial and often dominant role in shaping the electron spectral function [2]. The electron-phonon (-ph) interaction leads to the broadening and renormalization of the quasiparticle peak [3, 4, 5, 6, 7], photoemission kinks in metals [8, 9], and additional structures such as satellite peaks [10, 11, 12, 13, 14, 15]. Understanding and predicting these effects is the foundation of electron-phonon physics.
Ab initio calculation of the -ph self-energy and spectral functions is usually limited to perturbation theory. Standard approaches use the perturbative Fan–Migdal (FM) self-energy [16, 17], solving the Dyson equation [10] or applying the cumulant expansion [18, 19, 20, 21, 22, 15, 23, 24, 25, 26] for spectral functions. However, these approaches often fail, causing divergent quasiparticle dispersion [27, 28, 26] or unphysical spectral properties [26]. Notably, this problem is not limited to strongly coupled systems: the negative curvature in the quasiparticle dispersion of the Fröhlich model [27] and the divergent linewidth of piezoelectric materials [29] illustrate failures at all coupling strengths. Advanced methods incorporating higher-order -ph interactions [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44] are limited to model Hamiltonians. Non-perturbative methods using molecular dynamics [45, 46] or special-displacement supercells [47, 48] are difficult to converge [46, 49], and cannot capture non-adiabatic -ph coupling, crucial in polar materials [28, 50].
@@
A material of particular interest for non-perturbative -ph coupling is monolayer InSe. While known for high electron mobility [58, 59], its valence band features a flat, Mexican-hat-like dispersion [60, 61, 62, 63, 64, 65, 66], leading to strong -ph interaction. In particular, state-of-the-art one-shot calculations predicted a dynamical splitting of the valence band into two peaks in the spectral function with a gap larger than 150 meV [67]. The hole mobility of monolayer InSe has been studied using the Boltzmann transport equation (BTE) [68, 69, 70] and the bubble approximation with the one-shot spectral function [67]. Yet, given the strong renormalization, non-perturbative corrections are expected to be essential in monolayer InSe.
In this work, we present a non-perturbative ab initio calculation of the -ph spectral function and conductivity. Our approach, the self-consistent (sc) method [36, 37, 38, 39, 40, 41, 42, 43, 44], extends the FM self-energy to include the self-consistently renormalized spectral functions, analogously to the sc method for electron-electron interactions [71]. Using self-consistent spectral functions, we compute conductivity using the bubble approximation. We find that self-consistency significantly affects the spectral function and conductivity of monolayer InSe. This effect is present in any materials and can be explored with the sc method.
@@
where is the Brillouin zone volume, the -ph matrix element, and the electron and phonon energies, the Fermi–Dirac distribution, , and the Bose-Einstein distribution. We neglect the band-off-diagonal self-energy, which is a valid approximation if the bands do not hybridize [74], as in InSe. We refer to Eq. (1) as the -FM self-energy, corresponding to a one-loop diagram (see Fig. LABEL:supp-fig:feynman(a) of the SM [72]) with bare electron and phonon Green’s functions, and , as in the approximation [75]. In this work, we assume that is obtained from Kohn–Sham density functional theory and that and -ph coupling come from adiabatic density functional perturbation theory [76, 77, 2]. We neglect the renormalization of the phonon Green’s function since we focus on the intrinsic and low doping regime where the phonon renormalization is small [2]. The total self-energy also contains the Debye–Waller (DW) term [5, 4, 3]:
@@
The energy-dependent spectral function is computed by solving the Dyson equation:
@@
where is the Heaviside function and the cumulant function. We adopt the retarded cumulant method [81, 82, 24] as we deal with finite temperatures. The cumulant function is approximated via second-order perturbation theory [81, 21, 82]:
@@
To incorporate non-perturbative -ph effects, we adopt the sc method, where the self-energy is obtained by replacing the bare electron Green’s function in the approximation with the dressed one (see Fig. LABEL:supp-fig:feynman(b) of the SM [72]). The imaginary part of the sc self-energy is given by [56]
@@
We remark that it is advantageous to first compute the imaginary part and then use the Kramers–Kronig relation to obtain the real part. The relation between the real parts of and involves a -dependent energy integral [56]. Separating Eq. (8) and Eq. (9) enables performing the -integration for the imaginary part first, followed by a -independent energy integral.
@@
Here, is the one-shot FM self-energy [Eq. (1)] from states outside the active space, , evaluated at the bare eigenvalue . This term can be efficiently computed by solving the linear Sternheimer equation [79, 80] and interpolating using Wannier function perturbation theory (WFPT) [83]. The static corrections in the second line of Eq. (12) are always included during the self-consistent calculation of the sc self-energy, i.e., Eq. (12) replaces Eq. (9). (See Sec. LABEL:supp-sec:active_space of the SM [72] for additional details.)
In practice, we initialize the sc self-energy as a constant with meV. We do not use any artificial broadening during the sc iteration, and the choice of does not affect the converged sc self-energy. We iterate until the maximum change of is below meV. The sc method is implemented as an extension of our ElectronPhonon.jl package [84], which uses quantities computed from Quantum ESPRESSO [85], Wannier90 [86], and EPW [87, 88]. Computational details are provided in Sec. LABEL:supp-sec:comp_details of the SM [72].
The sc method, also known as the self-consistent Migdal, self-consistent first-Born, or the non-crossing approximation, has been applied to the Holstein [89, 90, 36, 37, 38, 40, 41, 42, 43, 44] and Anderson–Holstein models [39]. This non-perturbative method sums the infinite series of non-crossing rainbow diagrams (see Fig. LABEL:supp-fig:feynman(c) of the SM [72]), including those with an arbitrary number of -ph vertices. It is expected to be valid from weak to intermediate coupling but will fail in the strong coupling regime, as in the case of self-trapped small polarons [91]. State-of-the-art ab initio calculations [56, 57] used a one-shot calculation of Eqs. (8, 9) (without ) with the spectral function obtained from dynamical mean-field theory to compute the -ph self-energy. Here, we self-consistently solve Eqs. (8-12) to include non-perturbative -ph effects on the spectral function.
@@
Figures 1(a-d) show the valence band spectral functions of monolayer InSe at 100 K, compared to the angle-resolved photoemission spectroscopy (ARPES) photointensity [73] measured at the same temperature. The RS dispersion has several kinks with inverted curvature near the valence band maximum (VBM), an artifact also observed for the Fröhlich model [27]. The cumulant spectral function shows a single broad peak with no satellites and a prominent dip at . The one-shot method yields two dynamically split bands with a splitting of 170 meV, consistent with Ref. [67]. However, this large splitting vanishes under self-consistency. The sc spectral function shows a broadened quasiparticle peak with multiple satellites, with a distance between the peaks of 30–50 meV, comparable to the LO phonon energy of 27 meV. An analysis of the difference between and sc self-energies is provided in the End Matter (Fig. 4). Our sc result is consistent with the absence of a 150 meV scale gap in the ARPES photointensity [64, 73]. Although both the sc and cumulant spectral functions are consistent with the experimental observation within the experimental resolution, higher-resolution ARPES may observe the satellites predicted by our sc calculation.
@@
The corresponding mobility is with the carrier concentration. Equation (13) has been used to calculate the conductivity due to electron-electron interaction [54], -ph interaction [24, 55, 46], or both [56, 57]. Here, we investigate how the self-consistent renormalization of the spectral function affects the direct current (DC) and alternating current (AC) mobility of monolayer InSe.
@@
To study transport in the intrinsic limit, we set a small carrier concentration, placing the chemical potential within the band gap. In this nondegenerate regime, mobility should be independent of carrier concentration [92, 93]. However, as detailed in the End Matter, we find that both the and cumulant spectral functions yield strongly -dependent mobility (Fig. 5). We attribute this artifact to the slow decay of the spectral function above the band edge, which induces a large number of carriers on the tail of the spectral function. To address this problem, we developed a truncation scheme that removes the tail above a cutoff energy (see Sec. LABEL:supp-sec:decayof the SM [72]), which results in a -independent mobility. However, this scheme cannot be applied to the cumulant spectral function of InSe due to its large broadening. The sc method does not suffer from this problem and yields a -independent mobility.
@@
Figure 2(a) shows the intrinsic DC mobility of InSe. While the Boltzmann transport equation (BTE) and self-energy relaxation time approximation (SERTA) mobilities [94] are nearly identical and agree with the results from Refs. [69, 70], the bubble mobility with the sc spectral function is 3–4 times larger. This suggests that the quasiparticle approximation significantly underestimates the mobility. Using the one-shot spectral function instead of the sc one leads to an overestimation by more than an order of magnitude. These results can be understood in terms of the effective electron scattering rates shown in Fig. 2(b), which are inversely proportional to the mobility. The SERTA and BTE mobilities are linked to the the imaginary part of the self-energy at the bare band energy:
@@
For the spectral functions, we define the effective scattering rates using the self-energy at the renormalized quasiparticle energy:
@@
The renormalized quasiparticle energies, and , are the energies where the spectral functions and are maximal. Near the VBM, the BTE scattering rates exhibit a peak, while the sc result shows a smoother variation and smaller values due to the self-consistent broadening. For the one-shot self-energy, the scattering rates near the VBM are tiny because the renormalized valence bands lie within the DFT band gap where the imaginary part of the self-energy is small. This behavior is an artifact of the one-shot approximation.
@@
Finally, we show in Fig. 3 the AC mobility calculated using the and sc spectral functions. For the one-shot spectral function, we observe a peak around 150–250 meV, corresponding to the gap shown in Fig. 1(c), as in Ref. [67]. This feature persists with a larger artificial broadening (see Fig. LABEL:supp-fig:ac_sc_broadening of the SM [72]). However, with the self-consistent spectral function, the unphysical peaks are smoothed out, due to the reduced splitting and the broader Drude peak.
@@
The self-consistent framework developed in this work offers a versatile tool for analyzing the non-perturbative effect of -ph coupling with moderate computational effort, bridging the gap between one-shot approximations and numerically exact methods. While we have focused on the spectral functions and conductivity, this framework can be extended to other properties such as superconductivity and indirect absorption [95, 96]. Our work also lays the groundwork for applying diagrammatic many-body approaches to -ph systems. The sc method could be extended to explore contributions beyond the Fan–Migdal diagram. Notably, the development of a fully self-consistent scheme using the renormalized phonon Green’s function [36, 37, 97], the inclusion of the dynamical screening [98, 99, 100], and the consideration of -ph vertex correction [40, 42, 101] are promising avenues for future work. Combining the sc formalism with dynamical mean-field theory [102, 103] and its diagrammatic extensions [104] to study transport from the interplay of electron-electron [54] and -ph coupling [105] beyond the one-shot approximation [56, 57] is another interesting direction for future research.
The source code and data associated with this work are available on the Materials Cloud Archive [106].
@@ Acknowledgements.
We thank Neil R. Wilson and Vladimir I. Fal’ko for kindly providing the data for Ref. [73], and the anonymous referees who provided invaluable insights which led to an improved version of this work. S.P. is a Research Associate of the Fonds de la Recherche Scientifique - FNRS. This work was supported by the Fonds de la Recherche Scientifique - FNRS under Grants number T.0183.23 (PDR) and T.W011.23 (PDR-WEAVE). This publication was supported by the Walloon Region in the strategic axe FRFS-WEL-T. Computational resources have been provided by the EuroHPC JU award granting access to MareNostrum5 at Barcelona Supercomputing Center (BSC), Spain (Project ID: EHPC-EXT-2023E02-050), and by the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the FRS-FNRS under Grant No. 2.5020.11 and computational resources on Lucia, the Tier-1 supercomputer of the Walloon Region with infrastructure funded by the Walloon Region under the grant agreement n°1910247.References
@@
- Nery and Mauri [2022] J. P. Nery and F. Mauri, Non-perturbative Green’s function method to determine the electronic spectral function due to electron-phonon interactions: Application to a graphene model from weak to strong coupling, ArXiv220311289 Cond-Mat (2022).
@@
At 300 K, the sc self-energy reaches around 0.1 eV. While this value is smaller than the full bandwidth of the valence band of InSe (1.2 eV, see Fig. LABEL:supp-fig:InSe_itp of the SM [72]), it is larger than the width of the Mexican-hat feature at the valence band edge (0.07 eV). Thus, the self-energy leads to a significant renormalization near the valence band edge, qualitatively altering the spectral function.
@@
Tail of the spectral functions— Compared to the and cumulant spectral functions, the sc spectral functions show a clear qualitative difference in insulators and nondegenerate semiconductors: the former exhibits an unphysically long tail in the mid-gap region, which is absent in the latter. Figures 5(a, b) show the density of states (DOS)
@@
and the hole DOS (). In Fig. 5(a), the DOS from the and the cumulant methods show a slow decay (gray shaded curves). This is a result of the Lorentzian broadening of the bare Green’s function (see Sec. LABEL:supp-sec:decay of the SM [72] for a detailed derivation.) Importantly, reducing the artificial broadening reduces the amplitude of the tail, but does not change the scaling. When multiplied by the hole occupation factor, which increases exponentially with energy up to the chemical potential, the hole DOS continues to increase above the renormalized band edge, peaking at the chemical potential [Fig. 5(b)]. As a result, in nondegenerate semiconductors, the and cumulant methods incorrectly predict an unphysically large amount of carriers in the mid-gap region, where the spectral function is small. In contrast, the sc spectral function decays faster than (red shaded curve), and the hole DOS peaks at the renormalized band edge as expected.
@@
Doping dependence of the conductivity— Figure 5(c) shows the doping dependence of the hole mobility of monolayer InSe at . Surprisingly, the (green plus markers) and cumulant mobility (blue squares) show significant changes as a function of hole concentration. This result contrasts sharply with what is typically observed in nondegenerate semiconductors with dominant -ph scattering [92, 93]. We attribute this artifact to the unphysical tail in the spectral functions, which places most of the holes near the chemical potential, where the actual spectral function is small. Due to this strong doping dependence, the intrinsic (low doping limit) mobility cannot be properly calculated using the and cumulant approximations. To alleviate this problem, we developed a truncation scheme (see Eq. (LABEL:supp-eq:trunc_occ_p_doped) of the SM [72]) that removes the tail of the spectral function beyond the local minima in the hole DOS. Then, we obtain a doping-independent Bubble- mobility (green crosses). However, this truncation could not be applied to the cumulant spectral functions of InSe due to its large broadening.
@@
In contrast, the sc mobility (orange circles) as well as the BTE mobility (blue and red triangles) remain doping-independent in the nondegenerate regime (white background). Doping dependence is observed only in the metallic regime (gray background), where the chemical potential lies within the valence band. Therefore, it is highly desirable to use the sc method, rather than the or cumulant approximations, for studying transport in nondegenerate semiconductors.

### v3

# Nonperturbative self-consistent electron-phonon spectral functions and transport

@@ Abstract
Electron-phonon coupling often dominates the electron spectral functions and carrier transport properties. However, studies of this effect in real materials have largely relied on perturbative one-shot methods due to the lack of a first-principles theoretical and computational framework. Here, we present a self-consistent theory and implementation for the nonperturbative calculations of spectral functions and conductivity due to electron-phonon coupling. Applying this method to monolayer InSe, we demonstrate that self-consistency qualitatively affects the spectral function and transport properties compared to state-of-the-art one-shot calculations and allow one to reconcile calculations with angle-resolved photoemission experiments. The developed method can be widely applied to materials with dominant electron-phonon coupling at moderate computational cost.
Spectral functions are essential in many-body physics, linking theory and experiment. Experimentally, spectral functions can be extracted using high-resolution angle-resolved photoemission spectroscopy (ARPES) [1]. Theoretically, spectral functions can be calculated from Green’s functions, including many-body correlation via the self-energy. The coupling between electrons and phonons, the collective vibration of the lattice, plays a crucial and often dominant role in shaping the electron spectral function [2]. The electron-phonon (-ph) interaction leads to the broadening and renormalization of the quasiparticle peak [3, 4, 5, 6, 7], photoemission kinks in metals [8, 9], and additional structures such as satellite peaks [10, 11, 12, 13, 14, 15]. Understanding and predicting these effects is the foundation of electron-phonon physics.
Ab initio calculation of the -ph self-energy and spectral functions is usually limited to perturbation theory. Standard approaches use the perturbative Fan–Migdal (FM) self-energy [16, 17], solving the Dyson equation [10] or applying the cumulant expansion [18, 19, 20, 21, 22, 15, 23, 24, 25, 26] for spectral functions. However, these approaches often fail, causing divergent quasiparticle dispersion [27, 28, 26] or unphysical spectral properties [26]. Notably, this problem is not limited to strong e-ph coupling: the negative curvature in the quasiparticle dispersion of the Fröhlich model [27] and the divergent linewidth of piezoelectric materials [29] illustrate failures at all coupling strengths. Advanced methods incorporating higher-order -ph interactions [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44] are limited to model Hamiltonians. Nonperturbative methods using molecular dynamics [45, 46] or special-displacement supercells [47, 48] are difficult to converge [46, 49], and cannot capture nonadiabatic -ph coupling, crucial in polar materials [28, 50].
@@
A material of particular interest for nonperturbative -ph coupling is monolayer InSe. While known for high electron mobility [58, 59], its valence band features a flat, Mexican-hat-like dispersion [60, 61, 62, 63, 64, 65, 66], leading to strong -ph interaction. In particular, state-of-the-art one-shot calculations predicted a dynamical splitting of the valence band into two peaks in the spectral function with a gap larger than 150 meV [67]. The hole mobility of monolayer InSe has been studied using the Boltzmann transport equation (BTE) [68, 69, 70] and the bubble approximation with the one-shot spectral function [67]. Yet, given the strong renormalization, nonperturbative corrections are expected to be essential in monolayer InSe.
In this Letter, we present a nonperturbative ab initio calculation of the -ph spectral function and conductivity. Our approach, the self-consistent (sc) method [36, 37, 38, 39, 40, 41, 42, 43, 44], extends the FM self-energy to include the self-consistently renormalized spectral functions, analogously to the sc method for electron-electron interactions [71]. Using self-consistent spectral functions, we compute conductivity using the bubble approximation. We find that self-consistency significantly affects the spectral function and conductivity of monolayer InSe. This effect is present in any materials and can be explored with the sc method.
@@
where is the Brillouin zone volume, the -ph matrix element, and the electron and phonon energies, is a positive infinitesimal broadening, the Fermi–Dirac distribution, , and the Bose-Einstein distribution. We neglect the band off-diagonal self-energy, which is a valid approximation if the bands do not hybridize [74], as in InSe. We refer to Eq. (1) as the -FM self-energy, corresponding to a one-loop diagram (see Fig. LABEL:supp-fig:feynman(a) of Supplemental Material (SM) [72]) with bare electron and phonon Green’s functions, and , as in the approximation [75]. In this Letter, we assume that is obtained from Kohn–Sham density functional theory and that and -ph coupling come from adiabatic density functional perturbation theory [76, 77, 2]. We neglect the renormalization of the phonon Green’s function since we focus on the intrinsic and low doping regime where the phonon renormalization is small [2]. The total self-energy also contains the Debye–Waller (DW) term [5, 4, 3],
@@
The energy-dependent spectral function is computed by solving the Dyson equation
@@
where is the Heaviside function and the cumulant function. We adopt the retarded cumulant method [81, 82, 24] as we deal with finite temperatures. The cumulant function is approximated via second-order perturbation theory [81, 21, 82],
@@
To incorporate nonperturbative -ph effects, we adopt the sc method, where the self-energy is obtained by replacing the bare electron Green’s function in the approximation with the dressed one [see Fig. LABEL:supp-fig:feynman(b) in SM [72]]. The imaginary part of the sc self-energy is given by [56]
@@
We remark that it is advantageous to first compute the imaginary part and then use the Kramers–Kronig relation to obtain the real part. The relation between the real parts of and involves a -dependent energy integral [56]. Separating Eq. (8) and Eq. (9) enables performing the integration for the imaginary part first, followed by a -independent energy integral.
@@
Here, is the one-shot FM self-energy [Eq. (1)] from states outside the active space, , evaluated at the bare eigenvalue . This term can be efficiently computed by solving the linear Sternheimer equation [79, 80] and interpolating using Wannier function perturbation theory (WFPT) [83]. The static corrections in the second line of Eq. (12) are always included during the self-consistent calculation of the sc self-energy, i.e., Eq. (12) replaces Eq. (9). (See Sec. LABEL:supp-sec:active_space of SM [72] for additional details.)
In practice, we initialize the sc self-energy as a constant with meV. We do not use any artificial broadening during the sc iteration, and the choice of does not affect the converged sc self-energy. We iterate until the maximum change of is below meV. The sc method is implemented as an extension of our ElectronPhonon.jl package [84], which uses quantities computed from Quantum ESPRESSO [85], Wannier90 [86], and EPW [87, 88]. Computational details are provided in Sec. LABEL:supp-sec:comp_details of SM [72].
The sc method, also known as the self-consistent Migdal, self-consistent first-Born, or the noncrossing approximation, has been applied to the Holstein [89, 90, 36, 37, 38, 40, 41, 42, 43, 44] and Anderson–Holstein models [39]. This nonperturbative method sums the infinite series of noncrossing rainbow diagrams [see Fig. LABEL:supp-fig:feynman(c) of SM [72]], including those with an arbitrary number of -ph vertices. It is expected to be valid from weak to intermediate coupling but will fail in the strong coupling regime, as in the case of self-trapped small polarons [91]. State-of-the-art ab initio calculations [56, 57] used a one-shot calculation of Eqs. (8, 9) (without ) with the spectral function obtained from dynamical mean-field theory to compute the -ph self-energy. Here, we self-consistently solve Eqs. (8-12) to include nonperturbative -ph effects on the spectral function.
@@
Figures 1(a)-(d) show the valence band spectral functions of monolayer InSe at 100 K, compared to the ARPES photointensity [73] measured at the same temperature. The RS dispersion has several kinks with inverted curvature near the valence band maximum (VBM), an artifact also observed for the Fröhlich model [27]. The cumulant spectral function shows a single broad peak with no satellites and a prominent dip at . The one-shot method yields two dynamically split bands with a splitting of 170 meV, consistent with Ref. [67]. However, this large splitting vanishes under self-consistency. The sc spectral function shows a broadened quasiparticle peak with multiple satellites, with a distance between the peaks of 30–50 meV, comparable to the longitudinal optical (LO) phonon energy of 27 meV. An analysis of the difference between and sc self-energies is provided in the End Matter (Fig. 4). Our sc result is consistent with the absence of a 150 meV scale gap in the ARPES photointensity [64, 73]. Although both the sc and cumulant spectral functions are consistent with the experimental observation within the experimental resolution, higher-resolution ARPES may observe the satellites predicted by our sc calculation.
@@
The corresponding mobility is with the carrier concentration. Equation (13) has been used to calculate the conductivity due to electron-electron interaction [54], -ph interaction [24, 55, 46], or both [56, 57]. Here, we investigate how the self-consistent renormalization of the spectral function affects the dc and ac mobility of monolayer InSe.
@@
To study transport in the intrinsic limit, we set a small carrier concentration, placing the chemical potential within the band gap. In this nondegenerate regime, mobility should be independent of carrier concentration [92, 93]. However, as detailed in the End Matter, we find that both the and cumulant spectral functions yield strongly -dependent mobility (Fig. 5). We attribute this artifact to the slow decay of the spectral function above the band edge, which induces a large number of carriers on the tail of the spectral function. To address this problem, we developed a truncation scheme that removes the tail above a cutoff energy (see Sec. LABEL:supp-sec:decayof SM [72]), which results in a -independent mobility. However, this scheme cannot be applied to the cumulant spectral function of InSe due to its large broadening. The sc method does not suffer from this problem and yields an -independent mobility.
@@
Figure 2(a) shows the intrinsic dc mobility of InSe. While the BTE and self-energy relaxation time approximation (SERTA) mobilities [94] are nearly identical and agree with the results from Refs. [69, 70], the bubble mobility with the sc spectral function is 3 to 4 times larger. This suggests that the quasiparticle approximation significantly underestimates the mobility. Using the one-shot spectral function instead of the sc one leads to an overestimation by more than an order of magnitude, and the mobility strongly depends on the artificial broadening , varying inversely with [see Fig. LABEL:supp-fig:dc_sc_broadening of SM [72]]. These results can be understood in terms of the effective electron scattering rates shown in Fig. 2(b), which are inversely proportional to the mobility. The SERTA and BTE mobilities are linked to the the imaginary part of the self-energy at the bare band energy,
@@
For the spectral functions, we define the effective scattering rates using the self-energy at the renormalized quasiparticle energy,
@@
The renormalized quasiparticle energies, and , are the energies where the spectral functions and are maximal. Near the VBM, the BTE scattering rates exhibit a peak, while the sc result shows a smoother variation and smaller values due to the self-consistent broadening. For the one-shot self-energy, the scattering rates near the VBM are tiny because the renormalized valence bands lie within the bare band gap where the imaginary part of the self-energy is small. This behavior is an artifact of the one-shot approximation.
@@
Finally, we show in Fig. 3 the ac mobility calculated using the and sc spectral functions. For the one-shot spectral function, we observe a peak around 150–250 meV, corresponding to the gap shown in Fig. 1(c), as in Ref. [67]. This feature persists with a larger artificial broadening (see Fig. LABEL:supp-fig:ac_sc_broadening of SM [72]). However, with the self-consistent spectral function, the unphysical peaks are smoothed out due to the reduced splitting and the broader Drude peak.
@@
The self-consistent framework developed in this work offers a versatile tool for analyzing the nonperturbative effect of -ph coupling with moderate computational effort, bridging the gap between one-shot approximations and numerically exact methods. While we have focused on the spectral functions and conductivity, this framework can be extended to other properties such as superconductivity and indirect absorption [95, 96]. Our work also lays the groundwork for applying diagrammatic many-body approaches to -ph systems. The sc method could be extended to explore contributions beyond the Fan–Migdal diagram. Notably, the development of a fully self-consistent scheme using the renormalized phonon Green’s function [36, 37, 97], the inclusion of the dynamical screening [98, 99, 100], and the consideration of -ph vertex correction [40, 42, 101] is a promising avenue for future work. Combining the sc formalism with dynamical mean-field theory [102, 103] and its diagrammatic extensions [104] to study transport from the interplay of electron-electron [54] and -ph coupling [105] beyond the one-shot approximation [56, 57] is another interesting direction for future research.
@@ Acknowledgements.
We thank Neil R. Wilson and Vladimir I. Fal’ko for kindly providing the data for Ref. [73], and the anonymous referees who provided invaluable insights that led to an improved version of this work. S.P. is a Research Associate of the Fonds de la Recherche Scientifique - FNRS. This work was supported by the Fonds de la Recherche Scientifique - FNRS under Grants number T.0183.23 (PDR) and T.W011.23 (PDR-WEAVE). This publication was supported by the Walloon Region in the strategic axe FRFS-WEL-T. Computational resources have been provided by the EuroHPC JU award granting access to MareNostrum5 at Barcelona Supercomputing Center (BSC), Spain (Project ID: EHPC-EXT-2023E02-050), and by the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the FRS-FNRS under Grant No. 2.5020.11 and computational resources on Lucia, the Tier-1 supercomputer of the Walloon Region with infrastructure funded by the Walloon Region under the grant agreement n°1910247. Data availability—The data that support the findings of this article are openly available on the Materials Cloud Archive [106].References
@@
- Nery and Mauri [2022] J. P. Nery and F. Mauri, Nonperturbative green’s function method to determine the electronic spectral function due to electron-phonon interactions: Application to a graphene model from weak to strong coupling, Phys. Rev. B 105, 245120 (2022).
@@
At 300 K, the sc self-energy reaches around 0.1 eV. While this value is smaller than the full bandwidth of the valence band of InSe (1.2 eV; see Fig. LABEL:supp-fig:InSe_itp of SM [72]), it is larger than the width of the Mexican-hat feature at the valence band edge (0.07 eV). Thus, the self-energy leads to a significant renormalization near the valence band edge, qualitatively altering the spectral function.
@@
Tail of the spectral functions—Compared to the and cumulant spectral functions, the sc spectral functions show a clear qualitative difference in insulators and nondegenerate semiconductors: the former exhibits an unphysically long tail in the midgap region, which is absent in the latter. Figures 5(a, b) show the density of states (DOS)
@@
and the hole DOS (). In Fig. 5(a), the DOS from the and the cumulant methods shows a slow decay (gray shaded curves). This is a result of the Lorentzian broadening of the bare Green’s function (see Sec. LABEL:supp-sec:decay of SM [72] for a detailed derivation.) Importantly, reducing the artificial broadening reduces the amplitude of the tail but does not change the scaling. When multiplied by the hole occupation factor, which increases exponentially with energy up to the chemical potential, the hole DOS continues to increase above the renormalized band edge, peaking at the chemical potential [Fig. 5(b)]. As a result, in nondegenerate semiconductors, the and cumulant methods incorrectly predict an unphysically large amount of carriers in the midgap region, where the spectral function is small. In contrast, the sc spectral function decays faster than (red shaded curve), and the hole DOS peaks at the renormalized band edge as expected.
@@
Doping dependence of the conductivity—Figure 5(c) shows the doping dependence of the hole mobility of monolayer InSe at . Surprisingly, the (green plus markers) and cumulant mobility (blue squares) show significant changes as a function of hole concentration. This result contrasts sharply with what is typically observed in nondegenerate semiconductors with dominant -ph scattering [92, 93]. We attribute this artifact to the unphysical tail in the spectral functions, which places most of the holes near the chemical potential, where the actual spectral function is small. Because of this strong doping dependence, the intrinsic (low doping limit) mobility cannot be properly calculated using the and cumulant approximations. To alleviate this problem, we developed a truncation scheme [see Eq. (LABEL:supp-eq:trunc_occ_p_doped) of SM [72]] that removes the tail of the spectral function beyond the local minima in the hole DOS. Then, we obtain a doping-independent Bubble- mobility (green crosses). However, this truncation could not be applied to the cumulant spectral functions of InSe due to its large broadening.
@@
In contrast, the sc mobility (orange circles) as well as the BTE and SERTA mobility (blue and red triangles) remain doping-independent in the nondegenerate regime (white background). Doping dependence is observed only in the metallic regime (gray background), where the chemical potential lies within the valence band. Therefore, it is highly desirable to use the sc method, rather than the or cumulant approximations, for studying transport in nondegenerate semiconductors.